If y1 = 2x+2 and y2 = -x^2/2 are the solutions of the equation y= xy' +( y')^2/2 then are the constant multipliers c1 , c2 are arbitrary, also the solutions of the given D.E. ? Is the sum y1+ y2 a solution ? Justify your answer.
1
Expert's answer
2019-03-21T09:59:58-0400

Let


"y=c_1y_1+c_2y_2""y=c_1(2x+2)+c_2(-{x^2 \over 2})"

Then


"y'=2c_1-c_2x"

"c_1(2x+2)+c_2(-{x^2 \over 2})=x(2c_1x-c_2x)+{(2c_1-c_2x)^2 \over2}"

"({1\over 2}c_2-{1\over 2}c_2)x^2+(2c_1-2c_1+2c_1c_2)x+2c_1-2{c_1}^2=0"

"2c_1c_2x+2c_1-2{c_1}^2=0"

We see that constant multipliers c1 and c2 are not arbitrary. We have to satisfy


"c_1c_2=0, c_1-{c_1}^2=0"

Therefore, the sum y1+y2 is not the solution of the equation.